# Knockoffs for low dimensions: changing the nominal level post-hoc to gain power while controlling the FDR

### v1

@@ 4 Post-hoc $\alpha$ with derandomized knockoffs
17:return $R^{\textnormal{dph}}$, $\alpha^{\textnormal{dph}}$
Remark 3.
To get rid of the (arbitrary) parameter $\alpha^{\textnormal{ebh}}$ in Algorithm 2, one could set it to default values. For example, if $\alpha^{\textnormal{ebh}}=0$, then the derandomized ph-knockoff procedure chooses the smallest $\alpha^{\textnormal{dph}}\in(0,1]$ that leads to at least one rejection (and sets $\alpha^{\textnormal{dph}}=\alpha^{\textnormal{kn}}$ if no such $\alpha^{\textnormal{dph}}$ exists). Furthermore, if $\alpha^{\textnormal{ebh}}=1$, then Algorithm 2 chooses the smallest $\alpha^{\textnormal{dph}}\in(0,1]$ that leads to the same rejection set as RB-knockoff applied at level $\alpha^{\textnormal{ebh}}=1$.
Remark 4.
In case of $k=1$ the derandomized ph-knockoff method (Algorithm 2) and the ph-knockoff procedure (Algorithm 1) are equivalent regardless of the chosen $\alpha_{\textnormal{ebh}}\in[0,1]$. To see this, note that for $k=1$ we have that
Hence, if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$, then $R^{\textnormal{dph}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=R^{\textnormal{ph}}$ and $\alpha^{\textnormal{dph}}=p/(i^{\textnormal{ebh}}E_{(i^{\textnormal{ebh}})}^{\text{avg}})=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}}=\alpha^{\textnormal{ph}}$. Furthermore, if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\emptyset$, it holds that $i^{*}=\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}$, which implies $\alpha^{\textnormal{dph}}=\alpha^{\textnormal{ph}}$ and $R^{\textnormal{dph}}=R^{\textnormal{ph}}$ in the same manner as before. Hence, the post-hoc framework unifies the non-derandomized and the derandomized knockoff method.
@@ 4.1 A uniform improvement of the derandomized (ph-)knockoff procedure
In this paragraph, we introduce a uniform improvement of the derandomized knockoff procedure for fixed $\alpha^{\textnormal{ebh}}$ by Ren and Barber [21] and therefore also of the derandomized ph-knockoff procedure. However, this uniform improvement is based on the e-Closure Principle and we have not yet found a short-cut. This makes this uniform improvement computationally inefficient, which is why we still recommend Algorithm 2 for practical purposes.
@@
$E_{S}=\frac{1}{k}\left(\sum_{j=1}^{k}\frac{\sum_{i\in S}\mathbbm{1}\{W_{i}^{(j)}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}^{(j)}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}\right).$ (19)
Now, due to the proof for FDR control of BC-knockoff by Barber and Candès [1], it follows that one can also define local e-values by changing $[p]$ in the denominators to $S$:
@@
Obviously, $\tilde{E}_{S}\geq E_{S}$ for all $S\subseteq[p]$ and thus the e-Closure Principle with the local e-values $(\tilde{E}_{S})_{S\subseteq[p]}$, called closed knockoff in the following, never rejects less hypotheses than the RB-knockoff procedure. This was already noted by Xu et al. [36] for $k=1$ (knockoff without derandomization). However, while they showed that in case of $k=1$ closed knockoff is just equivalent to BC-knockoff, it becomes a real uniform improvement for $k>1$ as we illustrate in the following example.
Example 2.
@@ 4.2 Derandomized knockoffs for PFER control
Ren et al. [22] introduced a derandomized version of a knockoff procedure [12] for PFER control. We call this procedure RWC-knockoff in the following. The PFER is defined as the expected number of false discoveries $\textnormal{PFER}=\mathbb{E}[|R\cap I_{0}|]=\mathbb{E}[\textnormal{NFD}_{I_{0}}(R)]$, where $\textnormal{NFD}_{I_{0}}(R)=|R\cap I_{0}|$ for some rejection set $R$.
Let $\nu>0$ be an integer and $T_{\nu}^{\text{PFER}}=\inf\{t>0:\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=\nu-1\}$ be the stopping time for the PFER knockoff procedure introduced by Janson and Su [12]. Again, we denote by $T_{\nu}^{\text{PFER},(j)}$, $j\in[k]$, the stopping time in the $j$-th knockoff run. Then RWC-knockoff is given by the rejection set
@@
where $\eta\in(0,1]$ is some prespecified parameter that determines after how many “rejections” in the different knockoff runs a hypothesis is finally rejected. Ren et al. [22] showed that RWC-knockoff controls the PFER at level $\nu/\eta$. Again, there is no optimal choice for $\eta$ and they propose sample splitting to obtain a data-driven parameter. We show that $\eta$ can be chosen entirely based on the data while retaining a sensible guarantee.
Theorem 4.2.
@@
$\mathbb{E}\left[\sup_{\eta\in(0,1]}\textnormal{NFD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})\eta\right]\leq\nu.$ (22)
Hence, we can choose $\eta$ based on the data while maintaining $\mathbb{E}[\textnormal{NFD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})\eta]\leq\nu$.
Proof.
For the proof we implicitly use the e-Closure Principle for PFER control by Xu et al. [36]. Define the local e-values $E_{S}$, $S\subseteq[p]$, as in (18) but with the stopping times $T_{\nu}^{\text{PFER},(j)}$ instead of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$. To see that $\mathbb{E}_{H_{0}^{S}}[E_{S}]\leq 1$, see for example Remark 4 in [21]. For any $\eta\in(0,1]$ it holds that
$\frac{\eta}{\nu}\textnormal{NFD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})$ $=\frac{\eta}{\nu}\sum_{i\in I_{0}}\mathbbm{1}\left\{\left(\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right)\geq\eta\right\}$ (23)
$\leq\frac{1}{\nu}\sum_{i\in I_{0}}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}=E_{I_{0}},$ (24)
where the inequality follows from $\mathbbm{1}\{x\geq 1\}\leq x$ for all $x\geq 0$. Since $\mathbb{E}[E_{I_{0}}]\leq 1$, the claim follows. ∎
In this case we recommend to just calculate the rejection set $R_{\eta}^{\textnormal{RWC}}$ for each $\eta\in\{r/k:r\in[k]\}$ and then choose the combination of $\eta$ and $R_{\eta}^{\textnormal{RWC}}$ that fits best.
Remark 5.
In the same way as in Section 4.1 we can also uniformly improve RWC-knockoff by using $(\tilde{E}_{S})_{S\subseteq[p]}$ in (20) with $T_{\nu}^{\text{PFER},(j)}$ instead of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$ and then apply the e-Closure Principle for PFER control by Xu et al. [36].
@@ 5.1 Generating data
Our simulation framework builds on the setup of Ren and Barber [21]. As in the original work, we simulate data from both a Gaussian linear model and a logistic regression model. To reflect the two simulation regimes explored in Section 5.3—low-dimensional and high-dimensional—we describe each separately below.
@@
$\mathbf{X}\sim\mathcal{N}(0,\Sigma),\qquad\Sigma_{jk}=0.5^{|j-k|}.$ (25)
@@
$\textnormal{Gaussian model: }Y\mid\mathbf{X}\sim\mathcal{N}(\mathbf{X}^{\top}\beta,1),\qquad\textnormal{Logistic model: }Y\mid\mathbf{X}\sim\textnormal{Bernoulli}\!\left(\frac{e^{\mathbf{X}^{\top}\beta}}{1+e^{\mathbf{X}^{\top}\beta}}\right).$ (26)
@@
$\beta=\big(\underbrace{0,\ldots,0}_{z},\tfrac{\bar{\beta}_{1}}{\sqrt{n}},\underbrace{0,\ldots,0}_{z},-\tfrac{\bar{\beta}_{2}}{\sqrt{n}},\ldots\big).$ (27)
The spacing parameter $z$ is chosen as a function of $p$ and $p_{\text{relevant}}$, ensuring that signals are distributed evenly throughout the variable vector. As in Ren and Barber [21], this alternating-sign, evenly spaced construction reduces correlation between active variables while maintaining signal heterogeneity.
Low-dimensional regime (Section 5.3.1).
@@
$p=50,\qquad n=250,\qquad p_{\text{relevant}}\in\{1,\ldots,10\}.$ (28)
@@
$A\in\{2,\dots,10\}\text{ (Gaussian)},\qquad A\in\{8,\dots,16\}\text{ (Logistic)}.$ (29)
Results for this regime are presented in Figures 2 and 3, where we compare post-hoc knockoffs to the original knockoff procedure of Candès et al. [4].
High-dimensional regime (Section 5.3.2).
To enable direct comparison with the derandomized knockoff procedure of Ren and Barber [21], we also simulate $n=1000$ samples in the same high-dimensional settings, namely
$p=800\text{ (Gaussian)},\qquad p=600\text{ (Logistic)}.$ (30)
@@
$p_{\text{relevant}}\in\{20,40,80\}\text{ (Gaussian)},\qquad p_{\text{relevant}}\in\{20,40,60\}\text{ (Logistic)}.$ (31)
The signal amplitude ranges match those used in the low-dimensional study above. Results for this regime appear in Figures 4 and 5, where we benchmark our derandomized post-hoc procedure against the original derandomized knockoffs method.

### v2

@@ 4 Post-hoc $\alpha$ with derandomized knockoffs
17:return $R^{\textnormal{dph}}$, $\tilde{\alpha}^{\textnormal{dph}}$
4.1 Choosing the rejection set before the $\alpha$
Until now, we have still followed the classic order in multiple testing, where one first defines the $\alpha$ (in our case, possibly data-adaptive) and then determines the largest possible rejection set. In practice, it may well be interesting to reverse this. This means that, based on the data and contextual information like explainable relevance of a variable, we determine a rejection set $R$ that is of interest to us. Then, we choose the smallest $\alpha$ for which $R$ is contained in the closure $\mathcal{R}_{\alpha}(\boldsymbol{E})$ (see (2) for the definition of the closure set). Due to the simultaneous control of the e-Closure Principle over rejection sets and significance levels (3), this combination of $R$ and $\alpha$ provides post-hoc FDR control. In the following proposition we derive the smallest such $\alpha$ for the derandomized knockoffs above.
Proposition 6
Let $\boldsymbol{E}$ be the family of e-values defined by (9) and $R$ be a rejection set depending on the data in an arbitrary way. Then,
$\inf\left\{\alpha>0:R\in\mathcal{R}_{\alpha}(\boldsymbol{E})\right\}=\frac{p}{|R|\min_{i\in R}E_{i}^{\text{avg}}}\eqqcolon\tilde{\alpha}^{R}$
with the convention $\tilde{\alpha}^{R}=\infty$ if $\min_{i\in R}E_{i}^{\text{avg}}=0$. Therefore, we have
$\mathbb{E}\left[\frac{\textnormal{FDP}_{I_{0}}(R)}{\tilde{\alpha}^{R}}\right]\leq 1.$
Proof We want to find the smallest $\alpha>0$ such that $R\in\mathcal{R}_{\alpha}(\boldsymbol{E})$, which is equivalent to
$\frac{1}{p}\sum_{i\in S}E_{i}^{\text{avg}}\geq\frac{\textnormal{FDP}_{S}(R)}{\alpha}\text{ for all }S\subseteq[p].$
Since adding indices $i$ with $i\notin R$ to some $S$ keeps the right-hand side of the inequality the same but increases the left-hand side of the inequality, we can restrict to $S\subseteq R$. Let $E_{(1:R)}\geq\ldots\geq E_{(|R|:R)}$ be the sorted e-values with index in $R$. Then we can further reduce the property $R\in\mathcal{R}_{\alpha}(\boldsymbol{E})$ to
$\alpha\geq\frac{p\cdot s}{|R|\sum_{i=1+|R|-s}^{|R|}E_{(i:R)}}\text{ for all }s\in[|R|].$
Since the right-hand side of the inequality is maximized for $s=1$, the assertion follows.
In Section 6.2, we illustrate this idea of choosing the rejection set first based on a real data trial.
@@ 4.2 A uniform improvement of the derandomized (ph-)knockoff procedure
In this paragraph, we introduce a uniform improvement of the derandomized knockoff procedure for fixed $\alpha^{\textnormal{ebh}}$ by Ren and Barber (2024) and therefore also of the derandomized ph-knockoff procedure. However, this uniform improvement is based on the e-Closure Principle and we have not found a short-cut. This makes this uniform improvement computationally inefficient, which is why we still recommend Algorithm 2 for practical purposes.
@@
$E_{S}=\frac{1}{k}\left(\sum_{j=1}^{k}\frac{\sum_{i\in S}\mathbbm{1}\{W_{i}^{(j)}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}^{(j)}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}\right).$
Now, due to the proof for FDR control of BC-knockoff by Barber and Candès (2015), it follows that one can also define local e-values by changing $[p]$ in the denominators to $S$:
@@
Obviously, $\tilde{E}_{S}\geq E_{S}$ for all $S\subseteq[p]$ and thus the e-Closure Principle with the local e-values $(\tilde{E}_{S})_{S\subseteq[p]}$, called closed knockoff in the following, never rejects less hypotheses than the RB-knockoff procedure. This was already noted by Xu et al. (2025) for $k=1$ (knockoff without derandomization). However, while they showed that in case of $k=1$ closed knockoff is just equivalent to BC-knockoff, it becomes a real uniform improvement for $k>1$ as we illustrate in the following example.
Example 2
@@ 4.3 Derandomized knockoffs for PFER control
Ren et al. (2023) introduced a derandomized version of a knockoff procedure (Janson and Su, 2016) for PFER control. We call this procedure RWC-knockoff in the following. The PFER is defined as the expected number of false discoveries $\textnormal{PFER}=\mathbb{E}[|R\cap I_{0}|]=\mathbb{E}[\textnormal{FD}_{I_{0}}(R)]$, where $\textnormal{FD}_{I_{0}}(R)=|R\cap I_{0}|$ for some rejection set $R$.
Let $\nu>0$ be an integer and
$T_{\nu}^{\text{PFER}}=\inf\left\{t>0:\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=\nu-1\right\}$
be the stopping time for the PFER knockoff procedure introduced by Janson and Su (2016). Again, we denote by $T_{\nu}^{\text{PFER},(j)}$, $j\in[k]$, the stopping time in the $j$-th knockoff run. Then RWC-knockoff is given by the rejection set
@@
where $\eta\in(0,1]$ is some prespecified parameter that determines after how many “rejections” in the different knockoff runs a hypothesis is finally rejected. Ren et al. (2023) showed that RWC-knockoff controls the PFER at level $\nu/\eta$. Again, there is no optimal choice for $\eta$ and they propose sample splitting to obtain a data-driven parameter. We show that $\eta$ can be chosen entirely based on the data while retaining a sensible guarantee.
Theorem 7
@@
$\mathbb{E}\left[\sup_{\eta\in(0,1]}\textnormal{FD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})\eta\right]\leq\nu.$
Proof For the proof we implicitly use a modified version of the e-Closure Principle for PFER control by Xu et al. (2025). Define the local e-values $E_{S}$, $S\subseteq[p]$, as in (9) but with the stopping times $T_{\nu}^{\text{PFER},(j)}$ instead of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$. To see that $\mathbb{E}_{H_{0}^{S}}[E_{S}]\leq 1$, see for example Remark 4 in (Ren and Barber, 2024). For any $\eta\in(0,1]$ it holds that
$\frac{\eta}{\nu}\textnormal{FD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})$ $=\frac{\eta}{\nu}\sum_{i\in I_{0}}\mathbbm{1}\left\{\left(\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right)\geq\eta\right\}$
$\leq\frac{1}{\nu}\sum_{i\in I_{0}}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}=E_{I_{0}},$
where the inequality follows from $\mathbbm{1}\{x\geq 1\}\leq x$ for all $x\geq 0$. Since $\mathbb{E}[E_{I_{0}}]\leq 1$, the claim follows.
Hence, we can choose $\tilde{\eta}$ based on the data while maintaining
$\mathbb{E}\left[\frac{\textnormal{FD}_{I_{0}}(R_{\tilde{\eta}}^{\textnormal{RWC}})\tilde{\eta}}{\nu}\right]\leq 1.$ (12)
One possibility to choose $\tilde{\eta}$ in practice would be to just calculate $R_{\eta}^{\textnormal{RWC}}$ for every $\eta\in\{r/k:r\in[k]\}$ and then choose $\tilde{\eta}=\eta$ for the $\eta$ that leads to the most desirable $(R_{\eta}^{\textnormal{RWC}},\eta)$ pair. In Algorithm 3 we propose to choose the one that maximizes the product $R_{\eta}^{\textnormal{RWC}}\eta$. It should be noted that this implementation does not guarantee a “free-lunch” improvement over the original method, rejection set is not always larger. However, if the $(R_{\eta}^{\textnormal{RWC}},\eta)$ combination of the original method seems more desirable, we could also switch to that post-hoc.
Algorithm 3 Derandomized ph-knockoff for PFER control
Input: Knockoff level $\nu\in\mathbb{N}$ and knockoff statistics $W_{1}^{(j)},\ldots,W_{p}^{(j)}$ for each $j\in[k]$.
Output: Rejection set $R^{\textnormal{ph-}\nu}$ and post-hoc level $\tilde{\eta}^{\textnormal{ph-}\nu}$.
1:Define $\mathcal{K}=\{r/k:r\in[k]\}$
2:Define $\tilde{\eta}^{\textnormal{ph-}\nu}=\operatorname*{argmax}_{\eta\in\mathcal{K}}\{R_{\eta}^{\textnormal{RWC}}\eta\}$, where $R_{\eta}^{\textnormal{RWC}}$ is defined as in (11)
3:Set
$R^{\textnormal{ph-}\nu}=\left\{i\in[p]:\left(\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right)\geq\tilde{\eta}^{\textnormal{ph-}\nu}\right\}$
4:if $R^{\textnormal{ph-}\nu}=\emptyset$ then
5:  $\tilde{\eta}^{\textnormal{ph-}\nu}=1/2$
6:end if
7:return $R^{\textnormal{ph-}\nu}$, $\tilde{\eta}^{\textnormal{ph-}\nu}$
Remark 8
In the same way as in Section 4.2 we can also uniformly improve RWC-knockoff by using $(\tilde{E}_{S})_{S\subseteq[p]}$ in (10) with $T_{\nu}^{\text{PFER},(j)}$ instead of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$ and then apply the e-Closure Principle for PFER control by Xu et al. (2025).
Furthermore, as described in Section 4.1 for FDR control, we can also choose the rejection set first and then select the $\eta$ afterwards.
Proposition 9
Let $R$ be a rejection set depending on the data in an arbitrary way and
$\tilde{\eta}^{R}\coloneqq\min_{i\in R}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}.$
Then $\mathbb{E}[\textnormal{FD}_{I_{0}}(R)\tilde{\eta}^{R}]\leq\nu$.
Proof Let $E_{I_{0}}$ be defined as in the proof of Theorem 7. Then,
$\mathbb{E}[\textnormal{FD}_{I_{0}}(R)\tilde{\eta}^{R}/\nu]$ $=\mathbb{E}\left[\frac{1}{\nu}\sum_{i\in I_{0}}\mathbbm{1}\{i\in R\}\min_{l\in R}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{l}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right]$
$\leq\mathbb{E}\left[\frac{1}{\nu}\sum_{i\in I_{0}}\mathbbm{1}\{i\in R\}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right]$
$=\mathbb{E}\left[\frac{1}{\nu}\sum_{i\in I_{0}}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right]$
$=\mathbb{E}\left[E_{I_{0}}\right]\leq 1.$
@@ 5.1 Generating data
Our simulation framework builds on the setup of Ren and Barber (2024). As in the original work, we simulate data from both a Gaussian linear model and a logistic regression model. To reflect the two simulation regimes explored in Section 5.3, low-dimensional and high-dimensional, we describe each separately below.
@@
$\mathbf{X}\sim\mathcal{N}(0,\Sigma),\qquad\Sigma_{jk}=0.5^{|j-k|}.$
@@
$\textnormal{Gaussian model: }Y\mid\mathbf{X}\sim\mathcal{N}(\mathbf{X}^{\top}\beta,1),\qquad\textnormal{Logistic model: }Y\mid\mathbf{X}\sim\textnormal{Bernoulli}\!\left(\frac{e^{\mathbf{X}^{\top}\beta}}{1+e^{\mathbf{X}^{\top}\beta}}\right).$
@@
$\beta=\big(\underbrace{0,\ldots,0}_{z},\tfrac{\bar{\beta}_{1}}{\sqrt{n}},\underbrace{0,\ldots,0}_{z},-\tfrac{\bar{\beta}_{2}}{\sqrt{n}},\ldots\big).$
The spacing parameter $z$ is chosen as a function of $p$ and $p_{\text{relevant}}$, ensuring that signals are distributed evenly throughout the variable vector. As in Ren and Barber (2024), this alternating-sign, evenly spaced construction reduces correlation between active variables while maintaining signal heterogeneity.
5.1.1 Low-dimensional regime (Sections 5.3.1, 5.3.4 and 5.3.2)
@@
$p=50,\qquad n=250,\qquad p_{\text{relevant}}\in\{2,\ldots,10\}.$
@@
$A\in\{2,\dots,10\}\text{ (Gaussian)},\qquad A\in\{6,\dots,16\}\text{ (Logistic)}.$
Results for this regime are presented in:
  • Figures 2 and 3, where we compare our post-hoc knockoffs (Algorithm 1) to the original knockoff procedure of Candès et al. (2018).
  • Figure 4, where we compare our post-hoc knockoffs (Algorithm 1) to the calibration knockoffs procedure of Luo et al. (2025).
  • Figure 6, where we compare our derandomized post-hoc knockoffs for PFER control (Algorithm 3) against the original derandomized procedure for PFER proposed by Ren et al. (2023).
5.1.2 High-dimensional regime (Section 5.3.3)
To enable direct comparison with the derandomized knockoff procedure of Ren and Barber (2024), we also simulate $n=1000$ samples in the same high-dimensional settings, namely
$p=800\text{ (Gaussian)},\qquad p=600\text{ (Logistic)}.$
@@
$p_{\text{relevant}}\in\{20,80\}\text{ (Gaussian)},\qquad p_{\text{relevant}}\in\{15,60\}\text{ (Logistic)}.$
The signal amplitude ranges are $A\in\{2,\dots,8\}$ for the Gaussian and $A\in\{4,\dots,16\}$ for the Logistic models.
Results for this regime appear in Figure  5, where we benchmark our derandomized post-hoc procedure for controlling FDR (Algorithm 2) against the original derandomized knockoffs for FDR proposed by Ren and Barber (2024).